OpenStudy (adrianna.gongora):

Angie was working on solving the exponential equation 23x = 6; however, she is not quite sure where to start. Using complete sentences, describe to Angie how to solve this equation and how solving would be different if the bases were equal.

OpenStudy (adrianna.gongora):

23^x my bad

OpenStudy (wolf1728):

To solve an exponential equation take logarithms of both sides. Can you do that?

OpenStudy (adrianna.gongora):

could you give me an example? I think I know what you are talking about. c:

OpenStudy (wolf1728):

okay a good example is solve for x 23^x = 6 log (23^x) = log 6 x * log (23) = log(6)

OpenStudy (mathmusician):

you would have to take the natural log of both sides and move the exponent down using the logarithmic laws (i.e. xlog23 = log6).

OpenStudy (wolf1728):

MathMusician - you do not have to take natural logs of both sides. You can use base 10 logs or logs of any base.

OpenStudy (adrianna.gongora):

So I do that and solve it?

OpenStudy (adrianna.gongora):

x * log (23) = log(6) solve this I mean?

OpenStudy (wolf1728):

Well do you know how to calculate logs? A lot of calculators can. If not here's an online log calculator www.1728.org/logrithm.htm

OpenStudy (adrianna.gongora):

is there a base for this or no?

OpenStudy (wolf1728):

Clickable link: http://www.1728.org/logrithm.htm

OpenStudy (adrianna.gongora):

wait would I make 10 the base?

OpenStudy (wolf1728):

You can use any base but I like using base 10 as opposed to natural logs

OpenStudy (adrianna.gongora):

Okay I am gonna show you what I have and tell me if I did this right .

OpenStudy (adrianna.gongora):

x*log10(23)=log10(6) divide both sides by log10(23) x=log10(6)/log10(23) (decimal x=0.57144)

OpenStudy (wolf1728):

x * 1.361727836 = 0.77815125038 x = 0.77815125038 / 1.361727836 x = 0.5714440359 I'd say you did that correctly !!

OpenStudy (adrianna.gongora):

so would I put x*log10(23)=log10(6) divide both sides by log10(23) x=log10(6)/log10(23) (decimal x=0.57144) as my answer?

OpenStudy (wolf1728):

Yes, although I don't think you'd have to say decimal x = .57144 The answer is not a logarithm it is a regular number in base 10

OpenStudy (adrianna.gongora):

Okay thank you c: I have two more questions I am gonna post will you look at them?

OpenStudy (wolf1728):

well okay go ahead